I do not understand this..I have to write a system of 2 equations in 2 unknowns and solve each system by substitution.

thanks so so much

Answer by applesrbest(7)

Let Marysville by m and Springfield be s.
m+s = 25000 -------eqn 1
0.1m+0.09s = 2380 ----------eqn 2
From eqn 1: m = 25000-s ---------eqn 3
Sub eqn 3 into eqn 2:
(0.1)(25000-s)+0.09s = 2380
2500-0.01s = 2380
0.01s = 120
s = 12000
Sub s=12000 back into eqn 3:
m=13000